mikewwe13:

Winston is solving the equation 5x - 3(10 - 2x) = 9(2x + 6). The following shows the steps he has used so far to solve the equation. 5x - 3(10) - 3(-2x) = 9(2x) + 9(6) Step 2: 5x - 30 + 6x = 18x + 54 Which statement describes a possible next step that Winston can use to solve this equation? A. Add 6x to both sides of the equation. B. Subtract 30 from both sides of the equation. C. Add 18x to both sides of the equation. D. Subtract 5x from both sides of the equation.

mikewwe13:

@Vocaloid

Vocaloid:

there's only one step that will make the equation less complex than before and eliminate something try to apply each step (a,b,c,d) and see which one makes the equation simpler

mikewwe13:

i solved it earlier and i wanted to see the answer is C or A i'm debating

mikewwe13:

my pick is C

Vocaloid:

hint: try adding 18x to each side of the equation does the equation become simpler? is anything eliminated?

mikewwe13:

1. yes 2. no

Vocaloid:

5x - 30 + 6x = 18x + 54 5x - 30 + 6x + 18x = 18x + 18x + 54 notice how nothing can be eliminated? the equation is not any simpler than it was before, so C is not the answer try to slow down and start with each answer choice, actually doing the mathematics

mikewwe13:

ok now i know C issn't

Vocaloid:

please take each answer choice, step by step, and try doing the actual math before guessing

mikewwe13:

D.

Vocaloid:

good

mikewwe13:

cause the answer is -12

mikewwe13:

Which justification can be used for step 1 of the following equation. −3(2x+1) = 3x + 15 Step 1 −6x−3=3x+15 Step 2 −3=9x+15 Step 3 −18=9x Step 4 −2=x A. addition property of equality B. distributive property C. subtraction property of equality D. multiplication property of equality

Vocaloid:

hint - how did we get from −3(2x+1) to −6x−3

mikewwe13:

Multiplying using the distributive property

Vocaloid:

good, so B

mikewwe13:

Which justification can be used for step 2 of the following equation. -3(2x + 1) = 3x + 15 Step 1 −6x−3=3x+15 Step 2 −3=9x+15 Step 3 −18=9x Step 4 −2=x A. multiplication property of equality B. subtraction property of equality C. addition property of equality D. division property of equality

Vocaloid:

to go from step 1 to step 2 Step 1 −6x−3=3x+15 Step 2 −3=9x+15 we added 6x to both sides, giving us the addition property (C)